OpenStudy (anonymous):

log_10(6 x-1) - log_10(x-1) = 1

OpenStudy (anonymous):

Are you just putting up all your logarithm homework?

OpenStudy (anonymous):

no these are the ones i don't understand how to do

OpenStudy (anonymous):

log(x)-log(y)=log(x/y) as shown previously. 10^log_10(x/y)=10^whatever's on this side. Solve that way.

OpenStudy (agreene):

\[\log_{10}(6x-1)-\log_{10}(x-1)=1\] \[\large10^{\log_{10}(6x-1)}-10^{\log_{10}(x-1)}=10^1\] \[(6x-1)-(x-1)=10\] \[6x-1-x+1=10\] \[5x=10\] \[x=\frac{10}{5}\] \[x=2\]

OpenStudy (anonymous):

okay so whenever a multiplication you do subtraction? I thought that was for addition and subtraction was division.